# Normalizing Flows for Gaussian Process Regression under Hierarchical Shrinkage Priors

### v1

@@ 2.1 The Importance of Hierarchy
GPR models suffer from the curse of dimensionality in a specific way. Consider two points in the output space, $y$ and $y^{\prime}$. When a new covariate is added, the anisotropic distance in (4) can only increase [^1] Technically, it could remain constant if the associated $\theta_{j}=0$. However, given the use of continuous shrinkage priors, this will happen with probability 00., leading to reduced covariance between $y$ and $y^{\prime}$, and, more broadly, to an increasingly sparse covariance matrix. This sparsification occurs mechanically as the number of covariates grows, even if they do not influence the outcome of interest (Binois and Wycoff, 2022). This sparsification not only affects the covariance structure but can also lead to overfitting or reduced predictive accuracy, particularly in high-dimensional settings.
The priors on $\theta_{1},\dots,\theta_{d}$ are critical in mitigating the curse of dimensionality by pushing the $\theta_{j}$ values of irrelevant covariates toward zero, effectively excluding them. The choice of prior for the global shrinkage parameter $\tau$ plays a particularly important role here. As shown in Cadonna et al. (2020), under the triple gamma shrinkage prior, setting $\tau\sim F(2c,2a)$ implies a uniform prior on the model size, i.e. the number of relevant covariates. This ensures that some prior mass remains on models with few or no covariates, even when $d$ is large, while still allowing flexibility for larger models when supported by the data.
To be more precise on what ”model size” means, first consider the following equivalent re-parameterization of the triple gamma prior (Cadonna et al., 2020):
@@
This formulation generalizes the concept of shrinkage factors, well-studied in normal-means and regression settings (Carvalho et al., 2010), to positive-constrained parameters. The interpretation remains straightforward: as $\rho_{j}$ approaches zero, $\frac{\rho_{j}}{2(1-\rho_{j})}$ approaches zero, making the prior for $\theta_{j}$ converge to a gamma distribution with infinite mean and variance. Conversely, as $\rho_{j}$ approaches one, $\frac{\rho_{j}}{2(1-\rho_{j})}$ approaches infinity, driving the prior for $\theta_{j}$ toward a gamma distribution with mean and variance both zero. Thus, $\rho_{j}=1$ corresponds to total shrinkage, while $\rho_{j}=0$ implies no regularization on the parameter. Despite its origins in the sparse normal-means problem, this generalization of shrinkage factors adapts naturally to the square-transform case.
@@
What Cadonna et al. (2020) show is that under $\tau\sim F(2c,2a)$, the induced probability $\pi=\Pr(\rho_{j}<0.5)$ is uniformly distributed, implying a uniform prior on the model size. This is in stark contrast to a non-hierarchical prior, where $\pi$ is some fixed constant, a choice that is highly informative about the model size. The uniform prior ensures that models with few covariates are not overly penalized, which is critical for achieving robust performance in high-dimensional settings where irrelevant covariates dominate.
Examining the implied $\sum_{j=1}^{d}\theta_{j}$ can serve as a measure of the a priori distance between observations encoded within the Gaussian process prior. This is informative, as increased distance between obseravtions leads to a sparser covariance matrix. Figure 2 shows the distribution of this measure across various $d$ for both the hierarchical and non-hierarchical case. The hierarchical prior concentrates more mass on smaller values of $\sum_{j=1}^{d}\theta_{j}$, even as the covariate space increases. In contrast, the mode of the distribution under the non-hierarchical prior increases with the number of covariates, despite having the same conditional distribution. This behavior arises because the hierarchical prior induces a joint distribution over $\theta_{1},\dots,\theta_{d}$, where any large $\theta_{j}$ decreases the probability of other $\theta_{-j}$ values being large.
@@ 3 Estimation via Normalizing Flows
Estimation of GPR models is a notoriously difficult problem, despite the log likelihood (here marginalized over the function $f(\mathbf{x})$) being available in closed form:
@@
Two main problems arise: First, the triple gamma prior is not (conditionally or otherwise) conjugate to the posterior of the hyperparameters, which preculdes any closed form solutions and makes sampling based approaches less attractive. Second, the likelihood is expensive to evaluate, as it requires the inversion the $N\times N$ matrix $\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)+\sigma^{2}I$, which has a computational cost of order $\mathcal{O}(N^{3})$. This problem is compounded in the fully Bayesian approach, as sampling based methods often require the likelihood to be evaluated at each iteration. Many widely used variational methods, such as the mean-field approximation, alleviate the computational burden but often lack expressivity, leading to poor approximations of the posterior distribution.
This work proposes a different approach to the estimation of GPR models under hierarchical shrinkage priors. Specifically, it explores the use of normalizing flows (Rezende and Mohamed, 2015) to approximate the posterior distribution in a variational inference framework. While still an approximation to the true posterior, normalizing flows are highly flexible and can approximate even complex posterior shapes in a computationally tractable way. This makes them a step up from widely used mean field approximations, which, by design, are unable to capture correlations between parameters. This may be of particular importance in a regression setting, where explanatory variables are often correlated.
@@ 3.1 A Brief Introduction to Normalizing Flows
To illustrate, consider the goal of defining a joint distribution over a $D$-dimensional random variable $\mathbf{z}$ [^2] Note that, in this application, the $\mathbf{z}$ we wish to approximate corresponds to the posterior distribution. The key idea behind normalizing flows is to introduce a simple base distribution $p_{\mathbf{u}}(\mathbf{u})$ over a random variable $\mathbf{u}$ and then express $\mathbf{z}$ as a transformation $T$ of $\mathbf{u}$:
@@
Both the transformation $T$ and the base distribution $p_{\mathbf{u}}$ may depend on parameters, denoted here as $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$, respectively, leading to a distribution $p(\mathbf{z};\boldsymbol{\phi},\boldsymbol{\psi})$ over $\mathbf{z}$. For $T$ to qualify as a valid flow, it must be a diffeomorphism, meaning it is invertible and both $T$ and its inverse $T^{-1}$ are differentiable. This constraint also requires that $\mathbf{u}$ is of the same dimension $D$ as $\mathbf{z}$. Under these conditions, the density of $\mathbf{z}$ can be computed using the change of variables formula:
@@
where each $T_{k}$ is a transformation with parameters $\boldsymbol{\phi}_{k}$ and $\mathbf{u}_{k}$ represents the intermediate transformed variable after applying $T_{k}$. While each individual transformation $T_{k}$ may be simple, the composition of them is able to approximate more complex distributions. This structure is akin to a neural network for density estimation, with each transformation acting as a “layer” subject to constraints that ensure invertibility and differentiability.
@@ 3.1.1 An Example Transformation: The Radial Flow
There exist many possible transformations that could act as individual layers in the flow, such as the planar and radial flows from the seminal paper of Rezende and Mohamed (2015), the Inverse Autoregressive Flow (IAF) of Kingma et al. (2016), or the Sylvester normalizing flow of Berg et al. (2018), to name a few. While they differ in their specific form, a key property they all share is that they are designed such that the Jacobian and its determinant can be computed efficiently. This is mostly achieved through the use of triangular Jacobian matrices, which have a determinant that can be computed in $\mathcal{O}(D)$ time, where $D$ is the dimension of the random variable one wishes to model.
@@
with parameters $\beta,\alpha,\mathbf{u}_{0}$. The basic intuition behind this transformation is that it expands or contracts the base distribution around the point $\mathbf{u}_{0}$, either pushing mass towards $\mathbf{u}_{0}$ or pulling it away. The spread of the contraction is controlled by $\alpha$ and the direction and intensity is controlled by $\beta$. Through succesive applications of this transformation, it is possible to approximate distributions that are far more complex than would appear possible at first glance. Figure 3 shows the effect of radial flows on a 2D uniform distribution, with the base distribution progressively transformed by radial flows, introducing localized contractions and expansions. Note that each layer has its own parameters $\beta_{k},\alpha_{k},\mathbf{u}_{0,k}$, leading to different transformations at each step. These are the parameters that can be learned during estimation, with the goal of approximating the posterior distribution as closely as possible.
@@ 3.1.2 Normalizing Flow used in this Work
While the transformation behind the radial flow is simple and therefore good for building intuition, it suffers in terms of expressivity. As it can only contract or expand the base distribution around a single point, the number of layers $K$ needs to be large to make the flow expressive (Papamakarios et al., 2021). As such, more complex transformations are often used in practice, as in this work, where the Sylvester normalizing flow is employed (Berg et al., 2018):
$T(\mathbf{u})=\mathbf{u}+\mathbf{U}\mathbf{h}(\mathbf{W}^{T}\mathbf{u}+\mathbf{b}),$
where $\mathbf{U},\mathbf{W}\in\mathbb{R}^{D\times D}$, $\mathbf{u}\in\mathbb{R}^{D}$, $\mathbf{b}\in\mathbb{R}^{D}$ and $\mathbf{h}$ is an elementwise non-linear function (usually the hyperbolic tangent). It is a generalization of the planar flow of Rezende and Mohamed (2015) that increases the expressivity of each individual transformation. The increased expressivity of each individual layer means that fewer layers overall are required to reach the same level of expressivity.
Ensuring the transformation is invertible requires restricting the parameter space for $\mathbf{U}$, $\mathbf{W}$, and $\mathbf{b}$; see Berg et al. (2018) for details. It is important to note that the proposed methodology is not tied to Sylvester normalizing flows and can accommodate other transformations, if they are better suited to the problem at hand.
@@ 3.2 Normalizing Flows for Variational Inference
Normalizing flows’ capability to approximate complex distributions in a black-box fashion makes them well-suited for variational inference. In variational inference, the objective is to approximate the posterior distribution $p(\boldsymbol{\xi}\mid\mathbf{y})$ by selecting the best approximation from a family of distributions $q(\boldsymbol{\xi};\boldsymbol{\phi})$. This is achieved by minimizing the Kullback-Leibler (KL) divergence between the approximate distribution $q(\boldsymbol{\xi};\boldsymbol{\phi})$ and the true posterior $p(\boldsymbol{\xi}\mid\mathbf{y})$:
@@
where $\boldsymbol{\phi}$ now contains the parameters of each transformation $T_{k}$. Note that this is in contrast to the mean-field approximation, which assumes a factorized posterior with independent components, resulting in approximations that cannot capture dependencies. Normalizing flows, on the other hand, naturally capture dependencies within the posterior.
A common approach to minimizing the KL-divergence is by maximizing the Evidence Lower BOund (ELBO). When using normalizing flows as the approximate posterior, an ELBO of the following form arises (Papamakarios et al., 2021):
@@ 3.2.1 Practical Considerations
$-\lambda\sigma^{2},$ (17)
where (14) is the unnormalized version of the log likelihood in (3), (15) is a sum over unnormalized log densities of the triple gamma distribution in (9), (16) is the unnormalized log density of the F distribution, and (17) is the unnormalized log prior on $\sigma^{2}$, in this case assumed to be an exponential distribution.
From here, the gradients of the ELBO with respect to the parameters $\boldsymbol{\phi}$ can be computed via backpropagation, and the parameters can be updated using a gradient ascent method (e.g., through the pytorch library; Paszke et al., 2017). The entire procedure is summarized in Algorithm 1. The loop over the samples $S$ is written out for clarity, but, in practice, can be exectued in parallel, which makes the method highly efficient and well-suited for GPU computation.
@@ 4 Simulation Study
Figure 4: Log Predictive Density Scores (LPDS) for Maximum Likelihood (ML), mean-field (MF) approximation under a horseshoe (MF: HS) and triple gamma prior (MF: TG) and the proposed approach under a horseshoe prior (NF: HS) and triple gamma prior (NF: TG) across varying sparsity levels, sample sizes ($N$), and dimensions ($d$). NF methods outperform ML and MF, especially at higher sparsity and dimensionality, with NF: TG showing strong robustness, especially in high sparsity settings. Y-axes are scaled per row.
@@
  • Sparsity levels: The proportion of irrelevant covariates $s$, with $s\in\{0.5,0.6,0.7,0.8,0.9\}$.
  • Dimensionality: The number of covariates $d$, with $d\in\{5,10,25,50\}$.
  • Sample sizes: The number of observations $N$, with $N\in\{10,50,100,500\}$.
The covariates were generated as $x_{ij}\sim\mathcal{N}(0,1)$, while the inverse lengthscales $\theta_{1},\dots,\theta_{d}$ were drawn from a noncentral chi-squared distribution with 1 degree of freedom and a noncentrality parameter of $1.5^{2}$. This strikes a nice balance of ensuring the $\theta_{j}$ are different enough from 00 to be detected, while not having individual $\theta_{j}$ being so large they dominate the covariance structure. Of the $d$ $\theta_{j}$, $\lfloor s\cdot d\rfloor$ were randomly set to 00. The response variable was then generated as $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)+0.1\mathbf{I})$, where $\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)$ is the covariance matrix generated by the squared exponential kernel with parameters $\boldsymbol{\theta}$ and $\tau$.
The proposed normalizing flow approach was compared to a standard GPR model with an ARD kernel, estimated using a maximum likelihood (ML) approach with the GPy library (GPy, since 2012), as well as a mean-field variational approximation implemented in rstan (Stan Development Team, 2024). For the ML method, optimization was restarted 10 times, with the best result retained. For the mean-field approximation, convergence was assessed by ensuring that the relative norm of the objective was smaller than $0.01$, the default setting in rstan.
For both the proposed method and the mean-field approximation, two prior specifications were considered: one with $a=c=0.1$ and another with $a=c=0.5$. The former represents a more aggressive triple gamma specification, with strong shrinkage near the origin and very heavy tails, while the latter corresponds to the horseshoe prior of Carvalho et al. (2010). In both cases, $\tau\sim F(2c,2a)$ and $\sigma^{2}\sim\text{Exp}(10)$. The approximating flows consisted of 10 layers, and 10 latent samples were used for each Monte Carlo estimate of the ELBO. The normalizing flow models were run for 3000 iterations to ensure convergence, although this was typically achieved much earlier.

### v2

@@ 2.1 The Importance of Hierarchy
GPR models suffer from the curse of dimensionality in a specific way. Consider two points in the output space, $y$ and $y^{\prime}$. When a new covariate is added, the anisotropic distance in (3) can only increase [^1] Technically, it could remain constant if the associated $\theta_{j}=0$. However, given the use of continuous shrinkage priors, this will happen with probability 0., leading to reduced covariance between $y$ and $y^{\prime}$, and, more broadly, to an increasingly sparse covariance matrix. This sparsification occurs mechanically as the number of covariates grows, even if they do not influence the outcome of interest (binois2022survey). This sparsification not only affects the covariance structure but can also lead to overfitting or reduced predictive accuracy, particularly in high-dimensional settings.
The priors on $\theta_{1},\dots,\theta_{d}$ are critical in mitigating the curse of dimensionality by pushing the $\theta_{j}$ values of irrelevant covariates toward zero, effectively excluding them. The choice of prior for the global shrinkage parameter $\tau$ plays a particularly important role here. As shown in cad-etal:tri, under the triple gamma shrinkage prior, setting $\tau\sim F(2c,2a)$ implies a uniform prior on the model size, i.e., the number of relevant covariates. This ensures that some prior mass remains on models with few or no covariates, even when $d$ is large, while still allowing flexibility for larger models when supported by the data.
To be more precise on what "model size" means, first consider the following equivalent re-parameterization of the triple gamma prior (cad-etal:tri):
@@
This formulation generalizes the concept of shrinkage factors, well-studied in normal-means and regression settings (carvalho2010horseshoe), to positive-constrained parameters. The interpretation remains straightforward: as $\rho_{j}$ approaches zero, $\frac{\rho_{j}}{2(1-\rho_{j})}$ approaches zero, making the prior for $\theta_{j}$ converge to a gamma distribution with infinite mean and variance. Conversely, as $\rho_{j}$ approaches one, $\frac{\rho_{j}}{2(1-\rho_{j})}$ approaches infinity, driving the prior for $\theta_{j}$ toward a gamma distribution with mean and variance both zero. Thus, $\rho_{j}=1$ corresponds to total shrinkage, while $\rho_{j}=0$ implies no regularization on the parameter. Despite its origins in the sparse normal-means problem, this generalization of shrinkage factors adapts naturally to the square-transform case.
@@
What cad-etal:tri show is that under $\tau\sim F(2c,2a)$, the induced probability $\pi=\Pr(\rho_{j}<0.5)$ is uniformly distributed, implying a uniform prior on the model size. This is in stark contrast to a non-hierarchical prior, where $\pi$ is some fixed constant, a choice that is highly informative about the model size. The uniform prior ensures that models with few covariates are not overly penalized, which is critical for achieving robust performance in high-dimensional settings where irrelevant covariates dominate.
The implied $\sum_{j=1}^{d}\theta_{j}$ can serve as a measure of the a priori distance between observations encoded within the GP prior. Figure 2 shows the distribution of this measure across various $d$ for both the hierarchical and non-hierarchical triple gamma prior. The hierarchical prior concentrates more mass on smaller values of $\sum_{j=1}^{d}\theta_{j}$, even as the covariate space increases. In contrast, the mode of the distribution under the non-hierarchical prior increases with the number of covariates, despite having the same conditional distribution. This behavior arises because the hierarchical prior induces a joint distribution over $\theta_{1},\dots,\theta_{d}$, where any large $\theta_{j}$ decreases the probability of other $\theta_{-j}$ values being large.
@@ 3 Estimation via Normalizing Flows
Estimation of GPR models is a notoriously difficult problem, despite the log marginal likelihood (here marginalized over the function $f(\mathbf{x})$) being available in closed form:
@@
Two main problems arise: First, the triple gamma prior is not (conditionally or otherwise) conjugate to the posterior of the parameters, which precludes any closed form solutions and makes sampling based approaches less attractive. Second, the marginal likelihood is expensive to evaluate, as it requires the inversion the $N\times N$ matrix $\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)+\sigma^{2}I$, which has a computational cost of order $\mathcal{O}(N^{3})$. This problem is compounded in the fully Bayesian approach, as sampling based methods often require the marginal likelihood to be evaluated at each iteration. Many widely used variational methods, such as the mean-field approximation, alleviate the computational burden but often lack expressivity, leading to poor approximations of the posterior distribution.
This work proposes a different approach to the estimation of GPR models under hierarchical shrinkage priors. Specifically, it explores the use of normalizing flows (rezende2015variational) to approximate the posterior distribution in a variational inference framework. While still an approximation to the true posterior, normalizing flows are highly flexible and can approximate even complex posterior shapes in a computationally tractable way. This makes them a step up from widely used mean-field approximations, which, by design, are unable to capture correlations between parameters. This may be of particular importance in a regression setting, where explanatory variables are often correlated.
@@ 3.1 A Brief Introduction to Normalizing Flows
To illustrate, consider the goal of defining a joint distribution over a $D$-dimensional random variable $\mathbf{Z}$. [^2] Note that, in this application, the $\mathbf{Z}$ we wish to approximate is ultimately the posterior distribution. The key idea behind normalizing flows is to introduce a simple base distribution $p_{\mathbf{U}}(\mathbf{U})$ over a random variable $\mathbf{U}$ and then express $\mathbf{Z}$ as a transformation $T$ of $\mathbf{U}$:
@@
To be of interest for variational inference, the transformation $T$ depends on parameters, denoted here as $\boldsymbol{\phi}$, leading to a distribution $p_{\mathbf{Z}}(\mathbf{Z};\boldsymbol{\phi})$ over $\mathbf{Z}$ [^3] Technically, one could assume that the base distribution $p_{\mathbf{U}}(\mathbf{U})$ also depends on parameters. However, as will be seen later, one can often subsume the parameters of the base distribution into the transformation $T$, so they are omitted here.. Out of notational convenience, we will often drop the dependence on these parameters and write, e.g., $p_{\mathbf{Z}}(\mathbf{Z})$ instead. For $T$ to qualify as a valid flow, it must be a diffeomorphism, meaning it is invertible and both $T$ and its inverse $T^{-1}$ are differentiable. This constraint also requires that $\mathbf{U}$ is of the same dimension $D$ as $\mathbf{Z}$. Under these conditions, the density of $\mathbf{Z}$ can be computed using the change of variables formula:
@@
where each $T_{k}$ is a transformation with parameters $\boldsymbol{\phi}_{k}$ and $\mathbf{U}_{k}$ represents the intermediate transformed variable after applying $T_{k}$. While each individual transformation $T_{k}$ may be simple, the composition of them is able to approximate more complex distributions. This structure is akin to a neural network for density estimation, with each transformation acting as a "layer" subject to constraints that ensure invertibility and differentiability.
@@ 3.1.1 An Example Transformation: The Radial Flow
There exist many possible transformations that could act as individual layers in the flow, such as the planar and radial flows from the seminal paper of rezende2015variational, the Inverse Autoregressive Flow (IAF) of kingma2016improved, or the Sylvester normalizing flow of berg2018sylvester, to name a few. While they differ in their specific form, a key property they all share is that they are designed such that the Jacobian and its determinant can be computed efficiently. This is mostly achieved through the use of triangular Jacobian matrices, which have a determinant that can be computed in $\mathcal{O}(D)$ time, where $D$ is the dimension of the random variable one wishes to model.
@@
with parameters $\beta,\alpha,\mathbf{U}_{0}$. The basic intuition behind this transformation is that it expands or contracts the base distribution around the point $\mathbf{U}_{0}$, either pushing mass towards $\mathbf{U}_{0}$ or pulling it away. The spread of the contraction is controlled by $\alpha$ and the direction and intensity is controlled by $\beta$. Through successive applications of this transformation, it is possible to approximate distributions that are far more complex than would appear possible at first glance. Figure 3 shows the effect of radial flows on a 2D uniform distribution, with the base distribution progressively transformed by radial flows, introducing localized contractions and expansions. Note that each layer has its own parameters $\beta_{k},\alpha_{k},\mathbf{U}_{0,k}$, leading to different transformations at each step. These are the parameters that can be learned during estimation, with the goal of approximating the posterior distribution as closely as possible.
@@ 3.1.2 Normalizing Flow used in this Work
While the transformation behind the radial flow is simple and therefore good for building intuition, it suffers in terms of expressivity. As it can only contract or expand the base distribution around a single point, the number of layers $K$ needs to be large to make the flow expressive (papamakarios2021normalizing). As such, more complex transformations are often used in practice, as in this work, where the Sylvester normalizing flow is employed (berg2018sylvester):
$T(\mathbf{U};\mathbf{Q},\mathbf{W},\mathbf{b})=\mathbf{U}+\mathbf{Q}\mathbf{h}(\mathbf{W}^{T}\mathbf{U}+\mathbf{b}),$
where $\mathbf{U}\in\mathbb{R}^{D}$, $\mathbf{Q},\mathbf{W}\in\mathbb{R}^{D\times D}$, $\mathbf{b}\in\mathbb{R}^{D}$ and $\mathbf{h}$ is an elementwise non-linear function (usually the hyperbolic tangent). It is a generalization of the planar flow of rezende2015variational that increases the expressivity of each individual transformation. The increased expressivity of each individual layer means that fewer layers overall are required to reach the same level of overall expressivity.
Ensuring the transformation is invertible requires restricting the parameter space for $\mathbf{Q}$, $\mathbf{W}$, and $\mathbf{b}$; see berg2018sylvester for details. It is important to note that the proposed methodology is not tied to Sylvester normalizing flows and can accommodate other transformations, if they are better suited to a given problem at hand.
@@ 3.2 Normalizing Flows for Variational Inference
Normalizing flows’ capability to approximate complex distributions in a black-box fashion makes them well-suited for variational inference. If we let $\boldsymbol{\xi}$ contain all the parameters of a model we wish to estimate, then the objective of variational inference is to approximate the posterior distribution $p(\boldsymbol{\xi}\mid\mathbf{y})$ by selecting the best approximation from a family of distributions $q(\boldsymbol{\xi};\boldsymbol{\phi})$, which in turn depends on the parameters $\boldsymbol{\phi}$. This is achieved by minimizing the Kullback-Leibler (KL) divergence between the approximate distribution $q(\boldsymbol{\xi};\boldsymbol{\phi})$ and the true posterior $p(\boldsymbol{\xi}\mid\mathbf{y})$:
@@
where $\boldsymbol{\phi}$ now contains the parameters of all transformations $T_{1},\ldots,T_{K}$. Note that this is in contrast to the mean-field approximation, which assumes a factorized posterior with independent components, resulting in approximations that cannot capture dependencies. Normalizing flows, on the other hand, naturally capture dependencies within the posterior, as the transformations $T_{k}$ explicitly induce dependencies between the parameters.
A common approach to minimizing the KL-divergence is to maximize the Evidence Lower BOund (ELBO). When using normalizing flows as the approximate posterior, an ELBO of the following form arises (papamakarios2021normalizing):
@@ 3.2.1 Practical Considerations
$-\lambda\sigma^{2}$ (16)
$+C,$ (17)
where (13) is the unnormalized version of the log marginal likelihood in (3), (14) is a sum over unnormalized log densities of the triple gamma distribution in (8), (15) is the unnormalized log density of the F distribution, (16) is the unnormalized log prior on $\sigma^{2}$, in this case assumed to be an exponential distribution, and (17) is a constant that does not depend on the parameters.
From here, the gradients of the ELBO with respect to the parameters $\boldsymbol{\phi}$ can be computed via backpropagation, and the parameters can be updated using a gradient ascent method (e.g., through the pytorch library; paszke2017automatic). The entire procedure is summarized in Algorithm 1. The loop over the samples $S$ is written out for clarity, but, in practice, can be exectued in parallel, which makes the method highly efficient and well-suited for GPU computation.
Figure 4: Log of time in seconds to compute 100 draws or iterations (depending on the method) for Hamiltonian Monte Carlo (HMC), mean-field (MF) approximation, maximum likelihood (ML) estimation, and the proposed normalizing flow (NF) approach. The number of covariates $d$ is fixed at 25.
The computational efficiency of the proposed method is illustrated in Figure 4, which shows the time required to compute 100 iterations or draws (depending on the method) for Hamiltonian Monte Carlo (HMC), mean-field variational inference (MF), maximum likelihood estimation (ML), and the proposed normalizing flow approach (NF). The number of covariates is fixed at $d=25$. HMC and MF were implemented using the rstan package (rstan), ML was implemented using the GPy library (gpy2014), and the NF method was implemented using torch for R (torchR). All benchmarks were run on a laptop with an Intel Core i7-9750H CPU and an NVIDIA GeForce RTX 2060 GPU.
The results should not be interpreted as exact runtime comparisons, as each method is implemented in a different software framework with distinct language backends and optimization infrastructure. However, the plot provides a practical sense of the relative computational cost across approaches.
The NF method exhibits stable runtime with increasing $N$, requiring approximately 20 seconds across the entire range. While it is slower than ML and MF for small $N$, this advantage reverses at $N\approx 500$ for MF and at $N\approx 900$ for ML. This suggests that the main bottleneck in the NF implementation is likely due to fixed overhead in the underlying torch interface, rather than the complexity of the method itself. Once the bottleneck becomes the inversion of the covariance matrix, one would expect to see it scale similarly to the other methods. In contrast, HMC is the most computationally expensive method across all sample sizes, and becomes infeasible for larger datasets. As such, it was also excluded from the simulation study below, as it was not able to run in a reasonable time for larger datasets.
@@ 4 Simulation Study
Figure 5: Log Predictive Density Scores (LPDS) for Maximum Likelihood (ML), mean-field (MF) approximation under a horseshoe (MF: HS) and triple gamma prior (MF: TG) and the proposed approach under a horseshoe prior (NF: HS) and triple gamma prior (NF: TG) with multicollinearity level $\rho=0.5$ across varying sparsity levels, sample sizes (NN), and dimensions ($d$). Y-axes are scaled per row.
Figure 6: Log Predictive Density Scores (LPDS) for Maximum Likelihood (ML), mean-field (MF) approximation under a horseshoe (MF: HS) and triple gamma prior (MF: TG) and the proposed approach under a horseshoe prior (NF: HS) and triple gamma prior (NF: TG) with multicollinearity level $\rho=0.9$ across varying sparsity levels, sample sizes (NN), and dimensions ($d$). Y-axes are scaled per row.
@@
  • Sparsity levels: The proportion of irrelevant covariates $s$, with $s\in\{0.5,0.7,0.9\}$.
  • Dimensionality: The number of covariates $d$, with $d\in\{10,50,200\}$.
  • Sample sizes: The number of observations $N$, with $N\in\{100,500,2500\}$.
  • The degree of multicollinearity: The correlation between covariates, with $\rho\in\{0.5,0.9\}$.
The covariates were generated as $\boldsymbol{x}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{B})$, where $\boldsymbol{B}$ is a matrix with 1 on the main diagonal and all off-diagonal elements set to $\rho$. The inverse lengthscales $\theta_{1},\dots,\theta_{d}$ were drawn from a gamma distribution with shape 6 and scale 24. This strikes a nice balance of ensuring the $\theta_{j}$ are different enough from 0 to be detected, while not having individual $\theta_{j}$ being so large they dominate the covariance structure. Of the $d$ $\theta_{j}$, $\lfloor s\cdot d\rfloor$ were randomly set to 0. The response variable was then generated as $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)+0.1\mathbf{I})$, where $\mathbf{K}(\mathbf{x};\boldsymbol{\theta},\tau)$ is the covariance matrix generated by the squared exponential kernel with parameters $\boldsymbol{\theta}$ and $\tau$, which was set to 2 for all runs.
The proposed normalizing flow approach was compared to a standard GPR model with an ARD kernel, estimated using a maximum likelihood approach with the GPy library (gpy2014), as well as a mean-field variational approximation implemented in rstan (rstan). For the maximum likelihood method, optimization was restarted 10 times, with the best result retained. For the mean-field approximation, convergence was assessed by ensuring that the relative norm of the objective was smaller than $0.01$, the default setting in rstan.
For both the proposed method and the mean-field approximation, two prior specifications were considered: one with $a=c=0.1$ and another with $a=c=0.5$. The former represents a more aggressive triple gamma specification, with strong shrinkage near the origin and very heavy tails, while the latter corresponds to the variance shrinkage counterpart of the horseshoe prior of carvalho2010horseshoe. In both cases, $\tau\sim F(2c,2a)$ and $\sigma^{2}\sim\text{Exp}(10)$. The approximating flows consisted of 10 layers, and 10 latent samples were used for each Monte Carlo estimate of the ELBO. To save on memory, the number of latent samples was reduced to 2 in the case of $N=2500$. This did not seem to affect performance significantly. The normalizing flow models were run for 3000 iterations to ensure convergence, although this was typically achieved much earlier.